### v1

@@ H Results on Synthetic Dataset
We also conduct experiments in a synthetic video dataset. We evaluate the compared methods with paired clean videos that are spatially aligned with the smoky videos. We collect 380 clean video clips from public Cholec80 dataset [56], where 280 clips are used for training and the remaining 100 ones are used for evaluation. We follow the surgery smoke simulation manner [25] and add synthetic smoke in the clean video clips, except the first frame (regarded as PS frame). The results are shown in Tab. D. As the PS frame in the synthetic dataset is clean, we do not perform experiments with ${}^{\ast}$ and ${}^{\ast}$-S. Our results get the best PSNR scores, indicating the effectiveness of the proposed method. Moreover, the visual comparisons in Figs. J and K show that our results produce few visual artifacts, remove more clean smoke and are more consistent with the GT.
@@ I Network Details
Details of SelfSVD and ${}^{\ast}$. For SelfSVD and ${}^{\ast}$, the encoder includes two $3\times 3$ convolutional layers with a stride of 2 for scale down-sampling and 5 residual blocks [22] for feature extraction. We deploy individual encoders for the smoky frame and reference input respectively. The alignment module is built upon a pre-trained optical flow network (e.g., PWC-Net [54]) to achieve feature alignment. The fusion module consists of a $3\times 3$ convolutional layer for channel reduction and 60 residual blocks for feature enhancement. The reconstruction module includes 5 residual blocks, two pixel-shuffle operations, and a final $3\times 3$ convolutional layer.
@@
Figure J: Qualitative comparisons on the synthetic dataset. Our results produce few visual artifacts and are more consistent with the GT.
@@
Figure K: Qualitative comparisons on the synthetic dataset. Our results remove more clean smoke and are more consistent with the GT.
Figure L: Mask examples. A mask is generated to indicate areas with imperfect optical flow to prevent trivial solutions. Masked regions are marked with black pixels in (d).
@@
Layer
Filter
Stride
Output size
Conv, LeakyReLU
$3\rightarrow 64$
2
$128\times 128$
Conv, BN, LeakyReLU
$64\rightarrow 128$
2
$64\times 64$
Conv, BN, LeakyReLU
$128\rightarrow 256$
2
$32\times 32$
Conv, BN, LeakyReLU
$256\rightarrow 512$
2
$31\times 31$
Conv, BN, LeakyReLU
$512\rightarrow 1$
1
$30\times 30$
Details of SelfSVD-S and ${}^{\ast}$-S. The computation costs of SelfSVD and ${}^{\ast}$ are generally similar to the video processing methods. To make the computation cost consistent with single-image ones, we present the lightweight models SelfSVD-S and ${}^{\ast}$-S. Specifically, we first replace PWC-Net [54] in the alignment module to a more lightweight optical flow network SpyNet [46]. Second, we reduce the number of residual blocks in the feature encoder, masked-ref generator, fusion, and reconstruction module from 5, 5, 60, 5 to 3, 3, 8, 3 respectively. Third, we reduce the channel numbers from 64 to 32. Benefiting from the above simplification, SelfSVD-S and ${}^{\ast}$-S significantly reduce the number of model parameters and computation costs, while keeping performance.
@@
To prevent trivial solutions, we introduce a masking strategy, where a mask is generated to indicate areas with imperfect optical flow. Moreover, we process the smoky frame and warped reference (Ref) input with dark channel prior [21] (DCP) and high-degree Gaussian blur (Blur) to mitigate surgical smoke interference. Some mask examples are provided in Figs. L and M. On the one hand, as shown in Fig. L, it successfully detects the areas with imperfect optical flow. On the other hand, as shown in Fig. M, DCP and Blur help to avoid detecting incorrect masked areas.
Figure M: Effect of dark channel prior (DCP) and Gaussian blur (Blur) operations. DCP and Blur help to avoid detecting incorrect masked areas. The masked areas are marked with black pixels in (d), (e), and (f).
@@
Figure N: Effect of regularization term weight $\lambda_{reg}$. Please zoom in for better observation.
@@
$0.01$
24.46 / 0.6254 / 0.4650 / 4.91 / 3.88
$0.05$
24.23 / 0.6225 / 0.4626 / 4.85 / 3.87
$0.1$
24.03 / 0.6200 / 0.4658 / 4.87 / 3.82
K Effect of Regularization Term
Here we conduct experiments to validate the effect of regularization term weight $\lambda_{reg}$. The results are shown in Fig. N and Tab. F. In general, a smaller $\lambda_{reg}$ leads to a weaker suppression of trivial solutions, while a larger one leads to a weaker utilization of Ref. We set $\lambda_{reg}$ to 0.05 for the trade-off.
L Examples from LSVD Dataset
We construct the laparoscopic surgery video desmoking (LSVD) dataset from professional hospitals. Some examples from the dataset are provided in Fig. O. It can be seen that the dataset contains diverse and complex surgery smoke, such as mist, droplets, and streaks.
@@
In the main text, we utilize the processed PS frame as the target to evaluate desmoking methods, as smoke may remain in it. Here, we provide the quantitative comparison results when taking the original PS frame as the target, as shown in Tab. G. Due to interference from smoke in the PS frame, ${}^{\ast}$ and ${}^{\ast}$-S get lower PSNR and SSIM values than SelfSVD and SelfSVD-S, respectively. Moreover, it shows that the proposed methods still outperform state-of-the-art ones.
More qualitative comparison results are provided in Figs. P, Q, R and S. Our methods can remove more smoke, recover more photo-realistic details, and produce results more consistent with PS frames than state-of-the-art ones.
Table G: Quantitative comparison on LSVD dataset. The original PS frame is taken as the clean target for calculating metrics. The best results in each category are marked in bold.
@@
(Ours) ${}^{\ast}$ 24.62 0.6548
Figure O: Examples from LSVD dataset. Each row shows several frames from a video clip. It can be seen that the dataset contains diverse and complex surgery smoke, such as mist, droplets, and streaks. Please zoom in for better observation.
@@
Specifically, we prefer to select $\mathbf{S}_{ref}$ from previous neighboring frames that are clearer than $\mathbf{S}_{i}$. As shown in Algorithm 1, we first deploy a Ref detector according to the residual between the current smoky input and the desmoking output. $\mathbf{S}_{ref}$ is updated when next clearer frame occurs. Then we feed $\mathbf{S}_{ref}$ and the smoky video clip $\{\mathbf{S}_{k}\}_{k=i}^{i+L}$ into a SelfSVD model, generating the clean results $\{\mathbf{\hat{I}}_{k}\}_{k=i}^{i+L}$. $L$ is the frame number of the current video clip and we set it to 5. Please see some visualization examples at the site.
Algorithm 1 Pseudo code about the practical deployment of SelfSVD in surgery
@@
Figure P: Qualitative comparisons on LSVD dataset. Our methods generate results more consistent with the PS frame. Please zoom in for better observation.
@@
Figure Q: Qualitative comparisons on LSVD dataset. Our methods remove more clean smoke and recover more realistic details. Please zoom in for better observation.
@@
Figure R: Qualitative comparisons on LSVD dataset. Our methods generate results with fewer artifacts and are more consistent with the PS frame. Please zoom in for better observation.
@@
Figure S: Qualitative comparisons on LSVD dataset. Our methods remove more clean smoke and recover more fine-scale details. Please zoom in for better observation.

### v2

@@ 7 Conclusion
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant No. 62371164 and No. U22B2035.
@@ Acknowledgement
Renlong Wu Zhilu Zhang() Shuohao Zhang Longfei Gou
Haobin Chen Lei Zhang Hao Chen () Wangmeng Zuo
@@ H Results on Synthetic Dataset
We also conduct experiments in a synthetic video dataset. We evaluate the compared methods with paired clean videos that are spatially aligned with the smoky videos. We collect 380 clean video clips from public Cholec80 dataset [56], where 280 clips are used for training and the remaining 100 ones are used for evaluation. We follow the surgery smoke simulation manner [25] and add synthetic smoke in the clean video clips, except the first frame (regarded as PS frame). The results are shown in Tab. D. As the PS frame in the synthetic dataset is clean, we do not perform experiments with SelfSVD and SelfSVD-S. Our results get the best PSNR scores, indicating the effectiveness of the proposed method. Moreover, the visual comparisons in Figs. I and J show that our results produce few visual artifacts, remove more clean smoke and are more consistent with the GT.
@@ I Network Details
Details of SelfSVD and SelfSVD. For SelfSVD and SelfSVD, the encoder includes two $3\times 3$ convolutional layers with a stride of 2 for scale down-sampling and 5 residual blocks [22] for feature extraction. We deploy individual encoders for the smoky frame and reference input respectively. The alignment module is built upon a pre-trained optical flow network (e.g., PWC-Net [54]). The fusion module consists of a $3\times 3$ convolutional layer for channel reduction and 60 residual blocks for feature enhancement. The reconstruction module includes 5 residual blocks, two pixel-shuffle operations, and a final $3\times 3$ convolutional layer.
@@
Figure I: Qualitative comparisons on the synthetic dataset. Our results produce few visual artifacts and are more consistent with the GT.
@@
Figure J: Qualitative comparisons on the synthetic dataset. Our results remove more clean smoke and are more consistent with the GT.
Figure K: Mask examples. A mask is generated to indicate areas with imperfect optical flow to prevent trivial solutions. Masked regions are marked with black pixels in (d).
@@
Layer Filter Stride Output size
Conv, LeakyReLU $3\rightarrow 64$ 2 $128\times 128$
Conv, BN, LeakyReLU $64\rightarrow 128$ 2 $64\times 64$
Conv, BN, LeakyReLU $128\rightarrow 256$ 2 $32\times 32$
Conv, BN, LeakyReLU $256\rightarrow 512$ 2 $31\times 31$
Conv, BN, LeakyReLU $512\rightarrow 1$ 1 $30\times 30$
Details of SelfSVD-S and SelfSVD-S. The computation costs of SelfSVD and SelfSVD are generally similar to the video processing methods. To make the computation cost consistent with single-image ones, we present the lightweight models SelfSVD-S and SelfSVD-S. Specifically, we first replace PWC-Net [54] in the alignment module to a more lightweight optical flow network SpyNet [46]. Second, we reduce the number of residual blocks in the feature encoder, masked-ref generator, fusion, and reconstruction module from 5, 5, 60, 5 to 3, 3, 8, 3 respectively. Third, we reduce the channel numbers from 64 to 32. Benefiting from the above simplification, SelfSVD-S and SelfSVD-S significantly reduce the number of model parameters and computation costs, while keeping performance.
@@
To prevent trivial solutions, we generate a mask to indicate areas with imperfect optical flow. Moreover, we process the smoky frame and warped reference (Ref) input with dark channel prior [21] (DCP) and large-kernel Gaussian blur (Blur) to mitigate smoke interference. Mask examples are provided in Figs. K and L. On the one hand,, it successfully detects the areas with imperfect optical flow, as shown in Fig. K. On the other hand, DCP and Blur help to avoid detecting incorrect masked areas, as shown in Fig. L.
K Effect of Regularization Term
Here we conduct experiments to validate the effect of regularization term weight $\lambda_{reg}$. The results are shown in Fig. M and Tab. F. In general, a smaller $\lambda_{reg}$ leads to a weaker suppression of trivial solutions, while a larger one leads to a weaker utilization of Ref. We set $\lambda_{reg}$ to 0.05 for the trade-off.
L Examples from LSVD Dataset
We construct the laparoscopic surgery video desmoking (LSVD) dataset from professional hospitals. Some examples from the dataset are provided in Fig. N. It can be seen that the dataset contains diverse and complex surgery smoke, such as mist, droplets, and streaks.
Figure L: Effect of dark channel prior (DCP) and Gaussian blur (Blur) operations. DCP and Blur help to avoid detecting incorrect masked areas. The masked areas are marked with black pixels in (d), (e), and (f).
@@
Figure M: Effect of regularization term weight $\lambda_{reg}$. Please zoom in for better observation.
@@
$0.01$ 24.46 / 0.6254 / 0.4650 / 4.91 / 3.88
$0.05$ 24.23 / 0.6225 / 0.4626 / 4.85 / 3.87
$0.1$ 24.03 / 0.6200 / 0.4658 / 4.87 / 3.82
@@
In the main text, we utilize the processed PS frame as the target to evaluate desmoking methods, as smoke may remain in it. Here, we provide the comparison results when taking the original PS frame as the target, as shown in Tab. G. Due to interference from smoke in the PS frame, SelfSVD and SelfSVD-S get lower PSNR and SSIM values than SelfSVD and SelfSVD-S, respectively. Moreover, it shows that the proposed methods still outperform state-of-the-art ones.
More qualitative comparison results are provided in Figs. O, P, Q and R. Our methods can remove more smoke, recover more photo-realistic details, and produce results more consistent with PS frames than state-of-the-art ones.
Table G: Comparisons on LSVD dataset. The original PS frame is taken as the clean target for calculating metrics. The best results in each category are marked in bold.
@@
(Ours) SelfSVD 24.62 0.6548
Figure N: Examples from LSVD dataset. Each row shows several frames from a video clip. It can be seen that the dataset contains diverse and complex surgery smoke, such as mist, droplets, and streaks. Please zoom in for better observation.
@@
Specifically, we prefer to select $\mathbf{S}_{ref}$ from previous neighboring frames that are clearer than $\mathbf{S}_{i}$. As shown in Algorithm A, we first deploy a Ref detector according to the residual between the current smoky input and the desmoking output. $\mathbf{S}_{ref}$ is updated when next clearer frame occurs. Then we feed $\mathbf{S}_{ref}$ and the smoky video clip $\{\mathbf{S}_{k}\}_{k=i}^{i+L}$ into a SelfSVD model, generating the clean results $\{\mathbf{\hat{I}}_{k}\}_{k=i}^{i+L}$. $L$ is the frame number of the current video clip and we set it to 5. Please see some visualization examples at the https://github.com/ZcsrenlongZ/SelfSVD.
Algorithm A Pseudo code about the practical deployment of SelfSVD in surgery
@@
Figure O: Qualitative comparisons on LSVD dataset. Our methods generate results more consistent with the PS frame. Please zoom in for better observation.
@@
Figure P: Qualitative comparisons on LSVD dataset. Our methods remove more clean smoke and recover more realistic details. Please zoom in for better observation.
@@
Figure Q: Qualitative comparisons on LSVD dataset. Our methods generate results with fewer artifacts and are more consistent with the PS frame. Please zoom in for better observation.
@@
Figure R: Qualitative comparisons on LSVD dataset. Our methods remove more clean smoke and recover more fine-scale details. Please zoom in for better observation.